OpenStudy (anonymous):

Find the slope x- and y-intercepts of the graph of each equation. x-8y=3. I don't understand...

OpenStudy (anonymous):

First, let's arrange the equation into slope intercept form\[y = mx + b\]\(m\) is the slope. To find the x-int, set y=0. To find the y-int, set x=0.

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (anonymous):

Let me know what you get so I can double check your answer.

OpenStudy (anonymous):

Ok, just a prob... It says to find the slope, I already know it is the rise over run, but What do I do if I don't have a rise or a run?

OpenStudy (anonymous):

Put the equation into slope-intercept form, which is\[y = m x + b\]Let's arrange the given equation. \[-8y = -x + 3 \rightarrow y = { 1 \over8} x - {3 \over 8}\]

OpenStudy (anonymous):

Thanks, man, But i gotta go to sleep. C ya later...*sigh*